# Phase transition like behaviors of Propagation of Passenger Stranding phenomena in Subway Networks

### 12 Jan 2024 v1

@@ A.1 Data Description and Processing
A.1.1 Data Description
(1) Automatic Fare Collection data In this study, the actual passenger demand dataset for the service system is obtained from the Automatic Fare Collection (AFC) system of Beijing Rail Transit. The dataset comprises passenger travel data collected by the AFC system and includes all travel information from May 10th to May 14th, 2021. A total of 6,734,988 passenger records are obtained. The data attributes include TXN DATE TIME, DEVICE LOCATION, CARD SERIAL NUMBER, TRIP ORIGIN LOCATION, ENTRY TIME, PRODUCT ISSUER ID, PRODUCT TYPE, PAYMENT VALUE, CARD LIFE CYCLE COUNT, RECONCILIATION DATE, SETTLEMENT DATE, SAM ID, DEVICE ID, SOURCE PARTICIPANT ID, PURSE REMAINING VALUE, PURSE_REMAINING_VALUE, PTSN. The data used in this study are TXN DATE TIME, DEVICE LOCATION, CARD SERIAL NUMBER, TRIP ORIGIN LOCATION, ENTRY TIME.
(2) Subway Train Timetable Data
      (a)Train Classifications: The train in Beijing are categorized as A, B, C, D, and L. In the network model, all train on the lines are of type B, with a maximum capacity of 250 passengers per carriage. In the deal system: The passenger capacity of each train in the model is set at 1500 passengers per train. In the realistic model: The train models on each line in the network model are identical, but the train compositions are not entirely consistent on different lines. The maximum passenger capacity is solely determined by the train model and composition, resulting in slight variations in the maximum capacity among different lines. The train carriage models, compositions, and maximum passenger capacities for each line are listed in Table A.1. The actual capacity of the train carriages in the simulation system needs to be adjusted through parameter tuning, and the tuning results are presented in the following sections.
Table A.1: Train Type, Composition, and Maximum Passenger Capacity Information
Line Composition Maximum passenger capacity
@@ A.1.1 Data Description
(b) Categorized by Travel Segments: Shuttle trains and non-shuttle trains. The realistic system only includes shuttle trains, while the ideal system encompasses both shuttle and non-shuttle trains.
(3) Train Timetable Data The train timetable data represents the actual data for trains on the respective day, with the number of train services varying across different lines as shown in Table A.2. The Beijing subway timetable data used in this study primarily comes from the official application of Beijing Rail Transit App. This official app, maintained and released by the Beijing subway operating authority, provides comprehensive subway timetable information. The timetables include essential information such as train departure and arrival times, station details, and train frequencies for each line of the Beijing subway system.
Table A.2: The total number of train services on different lines on May 10th, 2021
@@
A.1.2 Data Processing
(1) AFC Data
      In this study, the network model consists of 9 out of all the lines in Beijing subway network. To process the original dataset, the following steps were taken:
      (a) Removal of records with missing key information, such as missing entry time, exit time, entry location, or exit location.
      (b) Elimination of duplicate records for the same station at the same time, both for entry and exit.
      (c) Exclusion of records with identical passenger ID, entry swipe time, exit swipe time, entry location, and exit location.
      (d) Exclusion of transaction records generated before 4:30 AM or after 12:30 AM, corresponding to the operational hours of the urban rail transit system.
      (e) Selection of records with entry and exit locations limited to stations on Lines 1, 2, 4, 5, 6, 7, 8, 10, and 13.
After these processing steps, a total of 3,811,418 AFC data transaction records that meet the research requirements were obtained. (2) Train Travel Data The train timetable data used in this study was obtained through web scraping of real-time data from the Beijing Rail Transit App. The data covers the period from May 10th to May 14th, 2021. Based on the acquired train timetable data, the departure time for each train was extracted. Using the timing information for the arrival at two consecutive stations, the travel time between each pair of stations was calculated as the time interval. These calculated intervals represent the travel time between stations.
A.2 2 Types of Systems
We have constructed an ideal system and a realistic system, with each system consisting of three components: the operation of trains, the generation of passengers, and the trains’ service to passengers.
A.2.1 The operation of trains
@@
Once the departure time is reached, the system automatically generates new trains from each originating station, without considering the source of the trains. The departure time, originating station, and destination station are set as follows:
      (a) In the ideal system The originating and destination stations for the trains are obtained from actual data (Table A.3), with each line having two directions of travel. On linear lines, the originating and destination stations are located at the two ends of each line. On circular lines, the originating and destination stations are adjacent stations. The direction of departure is from the originating station towards the destination station. From the start of the simulation, trains on each line start departing from their respective originating stations and continue until the end of the simulation, with a departure interval set at 4 min.
Table A.3: Information of Line Types, Direction, Originating Stations, and Destination Stations
@@ 2 Universal Report Gucheng
2 Circle 1 Jishuitan Xizhimen
@@ 2 Zhuxingzhuang Yinghai
10 Circle 1 Bagou Huoqiying
@@ 2 Dongzhimen Xizhimen
(b) In the realistic system
      The departure time, originating stations, and destination stations for both shuttle and non-shuttle trains on each line are obtained from actual data.
@@
      Trains always travel on a single line, starting from the originating station and sequentially passing through each station between the originating station and the destination station until reaching the destination station.
      Trains travel between adjacent stations, maintaining a constant speed during the travel process. The travel time is determined by the actual distance between stations. The actual interval data between stations is set as follows. The departure and arrival times of trains at each station in both directions are obtained from actual data. The line segments between stations are abstracted as edges between nodes, and following the directions specified in Fig. A.10, each edge on each line is numbered sequentially as "1, 2, 3, …" until all segments on the line are numbered. Based on the timetable data, the time intervals ($T_{i}$) between all stations in both directions of each line are calculated. Only the intervals satisfying $Ti\in(0,1000]$ seconds are selected. This time interval includes two components: the travel time of trains on the edges ($T_{r}$,0) and the service time to passengers ($T_{s}$) (set as 30s). By subtracting the service time ($T_{s}$) from the total interval, the travel time ($T_{i,1}$ and $T_{r,2}$) of trains on each edge of the line in both directions is obtained. The average of travel times on corresponding edges of both directions is calculated to obtain the travel time ($T_{i}$) of trains on each edge in the simulation system (Fig. A.10). This process starts from the departure and continues until all stations on the line have been traversed. After completion, trains are removed from the system, and post-travel scheduling is not considered.
Figure A.10: Travel time between stations. The direction of lines in the table is arranged according to the direction from originating station to destination station (direction 1) on each line.
@@ A.2.2 The Generation of Passengers
(1) Setting of Passenger Generation Time, Originating Stations, and Destination Stations
      The data input for the system is as follows: the passenger’s entry time serves as the time when the passenger is generated in the simulation, while the entry station and exit station represent the originating and destination stations in the system, respectively. The following describes the simulation of passengers in both the model system and the realistic model.
      (a) In the ideal model:
      At the start of the simulation, the system generates $n_{p}$ passengers every 1 min. The originating stations and destination stations for the generated passengers are randomly selected independently from $N$ stations ($N=234$). The passenger demand in the model system is denoted as $P_{u}$ and calculated as $n_{p}/N$.
      (b) In the realistic model:
      For the realistic passenger demand, the entry time and passenger OD are obtained from actual data, and the corresponding demand scale is set to 1.0.
      For additional passenger demands, $N_{d}$ instances are randomly selected from the original $N_{r}$ passenger data (each instance containing entry time, originating station, and destination station) to represent the modified entry passenger information, resulting in $N_{d}$ instances of passenger data. The adjusted passenger demand, denoted as $P_{r}$, is calculated as $N_{d}/N_{r}$.
      When the simulation time reaches the entry time of a passenger, the system generates the corresponding passenger at the originating station, and the passenger immediately enters the waiting state.
(2) Passenger Route Assignment
      We employ the Breadth-First Search (BFS) algorithm to find $k$ ($k=5$) shortest paths for each origin-destination (OD) pair. These paths are evaluated based on three factors: the number of stations passed ($n_{s}$), the total travel time ($T=T_{d}+T_{l}$, where $T_{d}$ represents the total travel time and $T_{l}$ represents the total station dwell time), and the number of transfer stations ($n_{t}$). A scoring mechanism is applied to the k paths, with the highest-scoring path selected as the route for that OD pair. The scoring rule is defined as follows:
$Score=(10/n_{s}+1000/T+1/n_{t})/111.$ (A.1)
@@ A.2.3 Trains’ Service to Passengers
(1) Service Process and Rules       Prior to the simulation, the number of trains, timetables for each train, and maximum passenger capacity per train have been remain unchanged throughout the simulation. For each arriving train, the following rules are followed in a specific order: first, passengers disembark before boarding; priority is given to passengers who arrived first, and if a train is already at maximum capacity, passengers cannot board. Specifically, the operations for trains stopped at stations are as follows:
      (a) The process of passengers getting off
      If the train reaches its final destination, all passengers onboard get off. Among the alighting passengers, those whose alighting station differs from their intended destination station need to transfer, otherwise they are removed. If the train has not reached its final destination, passengers who have reached their intended destination station need to get off and are removed. Among the remaining passengers who have not reached their destination, those whose next station on their path differs from the next station the train will arrive at need to transfer. The transfer rule is as follows: after alighting, passengers who need to transfer must queue for their next train on the platform that aligns with both the passenger and the train’s next station.
@@
      If a passenger’s next destination matches the train’s next station, they board the train in a first-come, first-served order. Otherwise, the passenger continues to wait in the queue.
@@
      The actual system duration is from 0 to 50400s and is collected every 5min.
      For the realistic system covered the period from 4:30 AM to 12:30 AM (a total of 73800s), data is collected every 10 min. (2) Dynamic Process of Passenger Boarding and Disembarkation The dynamic process of passenger boarding and alighting in the system can be illustrated by Fig.A.11, with specific symbols explained in Table A.4.
      We introduce $P^{\alpha}$ to represent the state of the train, and $P_{b}$ represents the station where the train is located. $P^{\alpha}=0$ indicates that the train has not yet departure, $P^{\alpha}=-1$ indicates that the train is loading passengers at the station, $P^{\alpha}=1$ represents the train is in travelling, and $P^{\alpha}=2$ indicates that the train has reached its destination. Passengers in the system are generated as follows. A passenger is generated every 1min, and their origin and destination are randomly selected from all stations. The number of generated passengers is set based on actual demand. Prior to the simulation, the passenger’s entry time, origin, and destination are set. During the simulation, the system utilizes the K-shortest path algorithm to determine the best travel route for passengers based on their origin and destination at the time of entry. The system records the travel demand status $S^{\alpha}$ for all generated passengers at each moment. There are four specific states: $S^{\alpha}=-1$ represents passengers waiting for boarding at the platform, $S^{\alpha}>0$ indicates that passengers have boarded and $S^{\alpha}$ represents the train number they are on, and $S^{\alpha}=0$ represents passengers have reached their destination.
Figure A.11: Passenger Boarding and Alighting Dynamics.
Table A.4: Notations, meaning and comments of the system
@@
$P^{\alpha}$ the state of trains $P^{\alpha}=-1$: loading passengers; $P^{\alpha}=1$: traveling; $P^{\alpha}=2$: arriving at destination.
$S^{\alpha}$ the state of passengers $S^{\alpha}=-1$: waiting; $S^{\alpha}>0$: being onboard, and $S^{\alpha}$ is the number of the train, $S^{\alpha}=0$: arriving at the destination.
$P^{s}$ the state of stations $P^{s}=1$: arriving at the destination; $P^{s}=2$: arriving at transfer stations, $P^{s}=0$: otherwise.
$P^{\beta}$ the location of the train current station number of the train
$P^{n}$ next station of the train next station number of the train
$P^{c}$ occupancy status of the train $P^{c}=1$: fully loaded; $P^{c}=0$: not fully loaded
$S^{d}$ passengers’ destination Passengers’ destination
$S^{n}$ next station of passengers next station of passengers
@@ A.2.4 Parameter settings and validation
Due to the fact that the network model does not include all the routes in Beijing’s entire transportation network, there are passengers in the real traffic flow whose origins or destinations are not located on the 9 lines of the network model. This leads to some deviations between the simulated system’s train capacity and the actual train capacity. To address this issue, we utilize the entry and exit timestamp from the AFC data to determine the travel time for passengers. By comparing the simulated travel time with the actual travel time derived from the AFC data, we adjust the parameter of train car capacity.
      Since different types of trains operate on the actual routes, there are differences in train capacities among different lines. The train capacities for each line are listed in Table A.1. During the parameter adjustment process, we maintain the proportional relationship between the car capacities of different lines consistent with the actual train capacities on each line. Let q0 represent the reference capacity of trains for Line 1, and q represents the adjusted train car capacity for Line 1, as shown in Table A.5.
Table A.5: Information on Reference Capacity and Design Standard Passenger Capacity of Subway Trains on Different Lines
Line Reference Capacity Standard Passenger Capacity
@@
      Step 1: We obtained the actual boarding time $t_{s,r}$ and alighting time $t_{e,r}$ for each passenger from the AFC data. We also obtained the corresponding passenger boarding time $t_{s,s}$($t_{s,r}=t_{s,s}$) and alighting time $t_{e,s}$ from the simulation system.
      Step 2: The actual travel time for each passenger is given by $\Delta t_{r}=t_{e,r}-t_{s,r}$, and the simulated travel time is given by $\Delta t_{s}=t_{e,s}-t_{s,s}$. The absolute error between the simulated and actual travel times is calculated as $\Delta t=\Delta t_{r}-\Delta t_{s}$.       Step 3: We applied the 3-sigma criterion to remove outliers from the set of absolute errors.
      Step 4: The absolute time errors are normalized to obtain the relative error for each passenger, defined as $\Delta t^{\prime}=\Delta t/\Delta t_{r}$. A smaller value of $\Delta t^{\prime}$ indicates that the simulation results for that passenger are closer to the actual data.
      Step 5: We calculated the mean $M_{t}$ and standard error ${SD}_{t}$ of the set of relative errors $\Delta t^{\prime}$. A mean value and variance closer to 0 indicate that the simulation results are closer to the actual data.
      By comparing the mean and variance of the relative errors for different train capacities (Fig. A.12), we observed that as the car capacity increases, both $M_{t}$ and $SD$ decrease. When $q=1.0$, $M_{t}$ and $S_{t}$ approach 0, indicating that the system corresponding to this set of car capacities is very close to the real system.
      In addition, we consulted the "Urban Rail Transit Engineering Project Construction Standard" (JB 104 2008), which specifies that in the case of overcrowding, the seating capacity remains unchanged, while the standing capacity is calculated as 9 persons/m2. Type B trains are comprised of 2 types of trains: single-driver trains and non-driver trains. The overcrowding threshold for single-driver trains is set at 230, while the overcrowding threshold for non-driver trains is set at 250. Each train set includes a single-driver train at each end, with the remaining cars being non-driver trains. The studied subway lines 1, 2, 4, 5, 6, 7, 8, 10, and 13 all use B-type trains, with the exception of lines 6 and 7, which have 8-car sets, while the rest of the lines have 6-car sets. Based on this information, we obtained the maximum passenger capacity for each line and train set (see Table A.5), which closely matches the parameter setting in our model, validating the effectiveness of our parameter adjustments.
Figure A.12: Variation of the Mean $M$ and Variance $SD$ of the Relative Error Set $\Delta t^{\prime}$ for the Simulated System with Different Train Capacities $q_{c}$. The blue curve represents the change in the mean $M$, and the orange curve represents the change in the variance $SD$. A closer approach to 0 for both $M$ and $SD$ indicates that the simulation model’s performance is closer to the actual data.

### 28 Feb 2024 v2

@@ A.1 Data description and processing
A.1.1 Passenger-related data
(1) Data description
In this study, the passenger demand dataset for the system II is obtained from the Automatic Fare Collection (AFC) system of Beijing Rail Transit. The dataset comprises passenger travel data collected by the AFC system and includes all travel information from May 10th to May 14th, 2021. The data attributes include TXN DATE TIME, DEVICE LOCATION, CARD SERIAL NUMBER, TRIP ORIGIN LOCATION, ENTRY TIME, PRODUCT ISSUER ID, PRODUCT TYPE, PAYMENT VALUE, CARD LIFE CYCLE COUNT, RECONCILIATION DATE, SETTLEMENT DATE, SAM ID, DEVICE ID, SOURCE PARTICIPANT ID, PURSE REMAINING VALUE, PURSE_REMAINING_VALUE, PTSN. The attributes of the data used in the study include TXN DATE TIME, DEVICE LOCATION, CARD SERIAL NUMBER, TRIP ORIGIN LOCATION, ENTRY TIME.
(2) Data processing
In this study, the network model comprises 9 out of the total 25 lines of the Beijing subway network as of December 2021. To process the original dataset, the following steps were taken:
(a) Remove records with missing TRIP ORIGIN LOCATION, ENTRY TIME, TXN DATE TIME and DEVICE LOCATION.
(b) Remove records which TRIP ORIGIN LOCATION and DEVICE LOCATION are the same.
(c) When multiple records have the same PRODUCT ISSUER ID, TRIP ORIGIN LOCATION, ENTRY TIME, TXN DATE TIME and DEVICE LOCATION as other records, keep only one of them.
(d) Remove records where the ENTRY TIME or TXN DATE TIME falls outside the operating hours of the subway (from 4:30 a.m. to 0:30 a.m. the next day).
(e) Remain records where both TRIP ORIGIN LOCATION and DEVICE LOCATION are limited to stations along Lines 1, 2, 4, 5, 6, 7, 8, 10, and 13.
A total of 15,850,040 passenger records were remained, including 3,811,418 records for Monday, 3,831,076 records for Tuesday, 3,833,882 records for Wednesday, 3,799,731 records for Thursday, and 4,023,933 records for Friday.
A.1.2 Subway-related data
(1) Data description
(a) Categorization of subway trains
Categorized by type, the trains in Beijing are classified into categories A, B, C, D, and L. In our system, all trains belong to type B, with a maximum capacity of 250 passengers per train. In system I, the passenger capacity of each train is set at 1500 passengers. In system II, the trains on each line are identical, but there may be slight variations in the maximum passenger capacity among different lines due to the train type and formation. The train types, compositions, and maximum passenger capacities for each line are listed in Table 1. The actual train capacity in system II needs to be adjusted through parameter tuning, and the tuning results will be presented in the following sections.
Table 1: Train Type, Formation, and Maximum Passenger Capacity Information
Line Formation Maximum passenger capacity
@@ A.1.2 Subway-related data
Categorized by shuttle interval, the trains in the system can be classified as shuttle trains and non-shuttle trains. System I exclusively consists of shuttle trains, while the composition of trains in System II is determined based on real data.
(b) Train timetable data
The train timetable in System II is set according to actual timetable. The number of train departures on each line are shown in Table 2. The train timetable data utilized in our study was obtained through web scraping of real-time data from the Beijing Rail Transit App. Beijing Rail Transit App is maintained and released by the Beijing subway operating authority, provides comprehensive subway timetable information. The timetables include essential details such as train departure and arrival times, station specifics, and train departure frequencies for each line in system II.
Table 2: The total number of train departures on each line
@@
(2) Data processing
Based on the acquired train timetable data, the departure time for each train was extracted. Using the timing information for the arrival at 2 adjacent stations, the travel time between each pair of stations was calculated as the time interval. These intervals represent the travel time between adjacent stations.
A.2 2 systems
We have established 2 systems: an ideal system and a realistic system, each comprising 3 components: train operation, passenger generation, and train service to passengers.
A.2.1 Train operation
@@
Once reached the departure time, both 2 systems automatically generate new trains from each originating station. The departure time, originating station, and destination station are set as follows:
(a) In system I
At the beginning of the simulation, trains on each line start departing from their respective originating stations and continue until the end of the simulation, with a departure interval set at 4 minutes. The originating and destination stations for the trains are obtained from real data (Table 3), and each line has 2 travel directions. For linear lines, the originating and destination stations are located at the 2 ends of each line. For circular lines, the originating and destination stations are adjacent stations. The departure direction is from the originating station towards the destination station.
Table 3: Information of Line Types, Direction, Originating Stations, and Destination Stations
@@ 2 Universal Report Gucheng
2 Loop 1 Jishuitan Xizhimen
@@ 2 Zhuxingzhuang Yinghai
10 Loop 1 Bagou Huoqiying
@@ 2 Dongzhimen Xizhimen
(b) In system II
The departure time, originating stations, and destination stations for both shuttle and non-shuttle trains on each line are obtained from real data.
@@
(a) System I
The trains consistently travel on a single line, departing from the originating station and sequentially passing through each station between the originating station and the destination station. Trains maintain a constant speed while traveling between adjacent stations. Since the distance between adjacent stations remains constant, the interval travel time can be determined using real data. The line segments between stations are represented as edges between nodes, and following the directions specified in Fig. 7, each edge on each line is sequentially numbered as “1, 2, 3, $\dots$” until all edges on the line are numbered. The average travel time taken by the trains to pass through is used as the interval travel time $T_{i}$ for the corresponding edges. Only intervals satisfying $T_{\mathrm{i}}\in(0,1000]$ seconds are retained. The interval travel time consists of 2 components: the travel time of trains ($T_{i}$,0) and the time spent at stations ($T_{\mathrm{s}}$) (set as 30s). By subtracting the dwell time ($T_{\mathrm{s}}$) from the total interval, the travel time ($T_{i\mathrm{,1}}$ and $T_{\mathrm{r,2}}$) of trains on each edge of the line in both directions is obtained. The average travel times on corresponding edges in both directions are calculated to determine the travel time ($T_{i}$) of trains on each edge in the simulation system (Fig. 7). This process starts from the departure and continues until all stations on the line would be traversed. Once the trains arrive at their destinations, they are removed from the system, and post-travel scheduling is not considered.
Figure 7: Travel time between stations. The direction of lines in the table is arranged according to the direction from originating station to destination station (direction 1) on each line.
@@ A.2.2 The generation of passengers
(1) Setting of passengers’ entry time, originating stations, exit time and destination stations
The attributes of the data inputted into the system includes: “TRIP ORIGIN LOCATION” ,“ENTRY TIME”,“ TXN DATE TIME” and “DEVICE LOCATION”, which respectively represent passengers’ originating stations, entry time, destination stations, and exit time. The following description outlines the simulation of passengers in both System I and System II.
(a) In system I
At the beginning of the simulation, System I generates $n_{\mathrm{p}}$ passengers every minute. The originating and destination stations for these passengers are randomly and independently selected from the entire pool of $N$ stations ($N=234$). The passenger demand in the system is denoted as $P_{\mathrm{u}}$ and calculated as $n_{\mathrm{p}}/N$.
(b) In system II
We randomly selected $N_{\mathrm{d}}$ records from the entire set of AFC data records, which comprised $N_{\mathrm{R}}$ records spanning from May 10th to May 14th, 2021. The selection was based on the records’ attributes of entry time and passenger origin-destination (OD) pairs. These $N_{\mathrm{d}}$ selected records were used to generate passengers in System II. Given that $N_{\mathrm{R}}$ represents the total records from 5 working days ($N_{\mathrm{R}}=19,300,040$), we defined the number of passengers per day as $N\mathrm{r}=N_{\mathrm{R}}/5$. The demand scale id defined as $P_{\mathrm{r}}=N_{\mathrm{d}}/N_{\mathrm{r}}$. When the number of selected passenger records, Nd, equaled $N_{\mathrm{r}}$, the demand scale of passengers, $P_{\mathrm{r}}$, would be 1.0. When the simulation time reaches the entry time of a passenger, the system generates the corresponding passenger at the originating station, and the passenger immediately transitions to the waiting state.
(2) Passenger route planning
The Breadth-First Search (BFS) algorithm is employed to identify the $k$ ($k=5$) shortest paths for each passenger’s OD pair. These paths are evaluated based on 3 factors: the number of stations passed ($n_{\mathrm{s}}$), the total travel time ($T=T_{\mathrm{d}}+T_{\mathrm{l}}$, where $T_{\mathrm{d}}$ represents the total travel time and $T_{\mathrm{l}}$ represents the total station dwell time), and the number of transfer stations ($n_{\mathrm{t}}$). A scoring mechanism is applied to the $k$ paths, and the highest-scoring path is selected as the route for that OD pair. The scoring rule is defined as follows:
$Score=(10/n_{\mathrm{s}}+1000/T+1/n_{\mathrm{t}})/111.$ (11)
@@ A.2.3 Trains’ service to passengers
(1) Service process and rules
Before the simulation begins, the number of trains, timetables for each train, and maximum passenger capacity per train are predetermined and remain consistent throughout the simulation. When a train arrives, the following rules are applied in a specific order: passengers disembark before boarding, priority is given to passengers based on their arrival time, and if a train has reached its maximum capacity, no additional passengers can board. Specifically, for trains stopped at stations:
(a) Passenger disembarkation
If the train reaches its destination, all passengers onboard should disembark. Among the disembarking passengers, those whose disembarkation station differs from their intended destination station need to transfer, while others are simply removed. If the train has not reached its destination, passengers who have reached their intended destination station need to get off and are removed. Among the remaining passengers who have not reached their destination, those whose next station on their path differs from the next station the train will arrive at need to transfer. The transfer rule is as follows: after disembarkation, passengers who need to transfer must queue for their next train on the platform that aligns with both the passenger and the train’s next station.
@@
If a passenger’s next station matches the train’s next station, they board the train in a first-come, first-served order. Otherwise, the passenger continues to wait in the queue.
@@
System I duration is from 0 to 50400s and is collected every 5min.
System II covered the period from 4:30 AM to 12:30 AM (a total of 73,800 seconds), with data collected every 10 minutes.
(2) Dynamic process of passenger boarding and disembarkation
The dynamic process of passenger boarding and disembarkation in the system can be illustrated by Fig. 8, with specific symbols explained in Table 4.
We introduce the variable $P^{\mathrm{\alpha}}$ to denote the state of the train, while $P_{\mathrm{b}}$ represents the station where the train is currently positioned. A value of $P^{\mathrm{\alpha}}=0$ indicates that the train has not yet departed, $P^{\mathrm{\alpha}}=-1$ signifies that the train is currently halted at a station, $P^{\mathrm{\alpha}}=1$ denotes that the train is in transit, and $P^{\mathrm{\alpha}}=2$ indicates that the train has reached its destination.
Passengers in the 2 systems are generated according to the following procedure. A passenger is generated every minute. Prior to the simulation, the passenger’s entry time, originating station, and destination station are predetermined. During the simulation, the system employs the K-shortest path algorithm to determine the optimal travel route for each passenger based on their starting and ending stations. The 2 systems track the travel demand status ($S^{\mathrm{\alpha}}$) of all generated passengers at each moment. There are 4 states: $S^{\mathrm{\alpha}}=-1$ represents passengers waiting to board at the platform, $S^{\mathrm{\alpha}}>0$ indicates that passengers have boarded and $S^{\mathrm{\alpha}}$ represents the train number they are on, and $S^{\mathrm{\alpha}}=0$ denotes passengers who have reached their destination station.
Figure 8: Passenger boarding and disembarkation dynamics.
Table 4: Notations, meaning and comments of the system
@@
$P^{\mathrm{\alpha}}$ Trains’ state $P^{\mathrm{\alpha}}=0$: not yet departed; $P^{\mathrm{\alpha}}=-1$: halted at the station; $P^{\mathrm{\alpha}}=1$: in transit; $P^{\mathrm{\alpha}}=2$: reached destination
$S^{\mathrm{\alpha}}$ Passengers’ state $S^{\mathrm{\alpha}}=-1$: waiting to board; $S^{\mathrm{\alpha}}>0$: being onboard, and $S^{\mathrm{\alpha}}$ train number they are on; $S^{\mathrm{\alpha}}=0$: reached destination station
$P^{\mathrm{s}}$ Stations’ type $P^{\mathrm{s}}=1$: reached destination; $P^{\mathrm{s}}=2$: reached at transfer station; $P^{\mathrm{s}}=0$: otherwise
$P^{\mathrm{\beta}}$ TrainS’ location Current station number of the train
$P^{\mathrm{n}}$ Next Station of the Train Next station number of the train
$P^{\mathrm{c}}$ Train occupancy state $P^{\mathrm{c}}=1$: fully loaded; $P^{\mathrm{c}}=0$: not fully loaded
$S^{\mathrm{d}}$ Passengers’ destination Passengers’ destination
$S^{\mathrm{n}}$ Next station of passengers Next station of passengers
@@ A.2.4 Parameter settings and validation
It is important to note that not all lines are included in the network model in this study. Consequently, there may be passengers within the real traffic flow whose origins or destinations are not situated along the 9 lines of the network model. This discrepancy gives rise to deviations between the train capacity simulated by the system and the actual train capacity. To tackle this challenge, we leverage the entry and exit timestamps obtained from the AFC data to determine the passengers’ travel time. By comparing the simulated travel time with the real travel time derived from the AFC data, we can make adjustments to the train capacity parameter, ensuring a more accurate representation of the system’s capacity.
As different types of trains on various lines, variations in train capacities arise across the network. The specific train capacities for each line are documented in Table 1. Throughout the parameter adjustment process, we ensure that the proportional relationship between train capacities on different lines remains consistent with the real train capacities observed for each line. In Table 5, we denote $q_{0}$ as the reference capacity of trains for Line 1, while $q$ represents the adjusted capacity of train cars for Line 1.
Table 5: Information on Reference Capacity and Design Standard Passenger Capacity of Subway Trains on Different Lines
Line Reference Capacity Standard Capacity
@@
Step 1: We obtained the real boarding time $t_{\mathrm{s,r}}$ and disembarkation time $t_{\mathrm{e,r}}$ for each passenger from the AFC data. We also obtained the corresponding passenger boarding time $t_{\mathrm{s,s}}$($t_{\mathrm{s,r}}=t_{\mathrm{s,s}}$) and disembarkation time $t_{\mathrm{e,s}}$ from the system II (set $P_{\mathrm{r}}=1.0$).
Step 2: The real travel time for each passenger is given by $\Delta t_{\mathrm{r}}=t_{\mathrm{e,r}}-t_{\mathrm{s,r}}$, and the simulated travel time is given by $\Delta t_{\mathrm{s}}=t_{\mathrm{e,s}}-t_{\mathrm{s,s}}$. The absolute error between the simulated and actual travel times is calculated as $\Delta t=\Delta t_{r}-\Delta t_{\mathrm{s}}$.
Step 3: We applied the 3-sigma criterion to remove outliers from the set of absolute errors.
Step 4: The absolute time errors are normalized to obtain the relative error for each passenger, defined as $\Delta t^{\prime}=\Delta t/\Delta t_{\mathrm{r}}$. A smaller value of $\Delta t^{\prime}$ indicates that the simulation results for that passenger are closer to the real data.
Step 5: We calculated the mean $M_{t}$ and standard error ${SD}_{t}$ of the set of relative errors $\Delta t^{\prime}$. A mean value and variance closer to 0 indicate that the simulation system are closer to the real-world system.
By comparing the mean and variance of the relative errors for different train capacities (Fig. 9), we observed that as the train capacity increases, both $M_{t}$ and $SD$ decrease. When $q=1.0$, $M_{t}$ and $S_{t}$ approach 0, indicating that the system corresponding to this set of train capacities is very close to the real-world system.
[…]
Figure 9: Mean ($M$) and variance ($SD$) of the relative error set ($\Delta t^{\prime}$) for the simulation system under different train capacities ($q_{\mathrm{c}}$). The blue curve represents the mean ($M$), while the orange curve represents the variance ($SD$).